On 29 March 2021, Justin deposited R3 500 into a savings account. The simple interest rate agreed upon was 7,5% per year. The accumulated amount in the savings account on 10 October 2021 is


[1] R3 640,24.

[2] R3 643,09.

[3] R3 637,88.

[4] R3 649,06.

[5] none of the above


1
Expert's answer
2021-05-17T18:52:31-0400

total days "=195\space days"

"amount=\frac{principal\times rate \times time}{100}"

"3500\times \frac{7.5}{100}\times{195}{365}"

"=140.24"

total accumulated amount"=3500+140.24=R\space 3640.24"

hence [A] is the right answer